[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jul 13, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jul 13 14:48:21 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:

HSU:
1. ATCF issues: Both for the 15Z and 21Z advisory, it took 3 tries sending  
the compute data for TC Eugene for the SDM to be able to see it.

These problems were likely caused by NCO turning off the LSF dev transfer  
queues on WCOSS because of network saturation (spiking to 99%) caused by  
data backfilling after the extended maintenance period required for  
splitting the GPFS filesystem on the Luna Cray supercomputing cluster.

The 24 hr parallel production test has ended, the data transfer catch-up  
process was completed at 6Z, and the LSF dev transfer queues have been  
reopened, so job submission and network traffic between NHC's ATCF servers  
and the WCOSS supercomputing clusters should return to normal.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170713/9a64124c/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list